# Scenario Analysis with Multivariate Bayesian Machine Learning Models

### v3

@@ B.1. Simulation-based results
Conditioning on shocks: GIRF computation. We use the same DGP as above, and treat the impact matrix $\bm{H}$ as well as the dynamic coefficients as given. This allows to compute the true impulse response function, which is shown in Figure B.2 as thick black line. We focus on the response of all variables to the first structural shock $j=1$. Panel (a) shows the distribution of predictive draws when using different random numbers for the scenario and baseline forecast, while panel (b) uses an identical sequence of random numbers for both forecasts in each iteration. In the latter case, there is no MCMC uncertainty (see also Appendix A), so our discussion below focuses on panel (a).

### v4

@@ B.1. Simulation-based results
Conditioning on shocks: GIRF computation. We use the same DGP as above, and treat the impact matrix $\bm{H}$ as well as the dynamic coefficients as given. This allows to compute the true impulse response function, which is shown in Figure B.2 as thick black line. We focus on the response of all variables to the first structural shock $j=1$. Panel (a) shows the distribution of predictive draws when using different random numbers for the scenario and baseline forecast, while panel (b) uses an identical sequence of random numbers for both forecasts in each iteration. In the latter case, there is no Monte Carlo uncertainty (see also Appendix A).